OpenStudy (anonymous):

While converting a triple integral from Cartesian to cylindrical coordinate system, how do I find the boundaries without drawing the graph?

OpenStudy (anonymous):

Okay well, it would help to have an example. But it takes a lot of algebraic manipulation.

OpenStudy (primeralph):

You'll need a good sense of imagination and a good understanding of functions; at least.

OpenStudy (anonymous):

You are basically manipulating inequalites

OpenStudy (anonymous):

\[\int\limits_{-1}^{1}\int\limits_{0}^{\sqrt{1-y ^{2}}}\int\limits_{0}^{x}(x^2+y^2)dzdxdy\]

OpenStudy (anonymous):

Okay so this is\[\begin{split} 0&<z&<x\\ 0&<x&<\sqrt{1-y^2}\\ -1&<y&<1 \end{split} \]

OpenStudy (anonymous):

Use your transformations on these inequalities and try to solve for \(r\) and \(\theta\)

OpenStudy (anonymous):

\[ \begin{split} 0&<z&<r\cos \theta \\ 0&<r\cos\theta &<\sqrt{1-r^2\sin^\theta }\\ -1&<r\sin \theta &<1 \end{split} \]

OpenStudy (anonymous):

The \(z\) bounds are already done.

OpenStudy (anonymous):

The second inequality can be squared \[ 0<r^2\cos^2 \theta <1-r^2\sin^2\theta \implies 0<r^2<1 \implies 0<r<1 \]

OpenStudy (anonymous):

Ah I see, great thanks, but what is wrong if I then transfer the inequalities to \[-1 \le y \le 1 , 0 \le x \le \sqrt{2} , 0 \le z \le \sqrt{2}\]

OpenStudy (anonymous):

You going to go from integrating over a nice rectangular cylinder to integrating over a screwed up region in cylindrical coords.

OpenStudy (anonymous):

The point is to make it easier, not harder.

OpenStudy (anonymous):

Is the approach entirely wrong?